OpenStudy (anonymous):

The time t required to drive a certain distance varies inversely with the speed r. If it takes 2 hours to drive the istance at 30miles per hour, how long will it take to drive the same distance at 50miles per hour ?

OpenStudy (a_clan):

t=2*30/50 =6/5 hrs

OpenStudy (anonymous):

2*30=60 miles - distance 60/50=1.2 hours